In a car with 4 cylinder, four stroke petrol engine having 120.6 mm bore and 132 mm stroke, the net dynamometer load is 195 N at a diameter of 1000 mm when the speed is 2500 rpm. At this speed and throttle opening the engine requires 5.2 kW to motor it with ignition switched off


i) Calculate the mechanical efficiency and the indicated mean effective pressure,


i) During the 5 minutes run at this speed and power, the engine used 1 kg of petrol of heating value 45310 kJ/kg and 40 kg of cooling water with a temperature rise of 55.5 K Specific heat of the water is taken as 4/18 kJ/kgK. Draw the heat balance chart of the test in kJ/min


1
Expert's answer
2021-07-05T08:43:30-0400
"MEP=\frac{4F}{\pi d^2}\\MEP=\frac{4(195)}{\pi (1)^2}=248\ Pa"